OpenStudy (anonymous):

(-7x - 5x^4 + 5) - (-7x^4 - 5 - 9x)

OpenStudy (apoorvk):

Subtract like terms. that is with like degrees of 'x'

OpenStudy (anonymous):

(-7x-5x^(4)+5)-(-7x^(4)-5-9x) Reorder the polynomial -7x-5x^(4)+5 alphabetically from left to right, starting with the highest order term. (-5x^(4)-7x+5)-(-7x^(4)-5-9x) Reorder the polynomial -7x^(4)-5-9x alphabetically from left to right, starting with the highest order term. (-5x^(4)-7x+5)-(-7x^(4)-9x-5) Remove the parentheses around the expression -5x^(4)-7x+5. -5x^(4)-7x+5-(-7x^(4)-9x-5) Multiply -1 by each term inside the parentheses (-7x^(4)-9x-5). -5x^(4)-7x+5+-(-7x^(4))-(-9x)-(-5) Multiply -1 by the -7x^(4) inside the parentheses. -5x^(4)-7x+5+7x^(4)-(-9x)-(-5) Multiply -1 by the -9x inside the parentheses. -5x^(4)-7x+5+7x^(4)+9x-(-5) Multiply -1 by the -5 inside the parentheses. -5x^(4)-7x+5+7x^(4)+9x+5 According to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. Here, x^(4) is a factor of both -5x^(4) and 7x^(4). (-5+7)x^(4)-7x+5+9x+5 To add integers with different signs, subtract their absolute values and give the result the same sign as the integer with the greater absolute value. In this example, subtract the absolute values of -5 and 7 and give the result the same sign as the integer with the greater absolute value. (2)x^(4)-7x+5+9x+5 Remove the parentheses. 2x^(4)-7x+5+9x+5 According to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. Here, x is a factor of both -7x and 9x. 2x^(4)+(-7+9)x+5+5 To add integers with different signs, subtract their absolute values and give the result the same sign as the integer with the greater absolute value. In this example, subtract the absolute values of -7 and 9 and give the result the same sign as the integer with the greater absolute value. 2x^(4)+(2)x+5+5 Remove the parentheses. 2x^(4)+2x+5+5 Add 5 to 5 to get 10. 2x^(4)+2x+10